# Nondeterminism-Aware Optimistic Verification for Floating-Point Neural Networks

### v2

@@ Abstract.
Neural networks increasingly run on hardware outside the user’s control (cloud GPUs, inference marketplaces, edge specialized accelerators) for both training and inference. Yet ML-as-a-Service reveals little about what actually ran or whether returned outputs faithfully reflect the intended inputs. Users lack recourse against service downgrades such as model swaps, quantization, graph rewrites, or discrepancies like altered advertisement embeddings. Verifying outputs is especially difficult because floating-point execution on heterogeneous accelerators is inherently nondeterministic. Existing approaches like zkML, deterministic replay, TEEs, and replication are either impractical for real floating-point neural networks or reintroduce vendor trust. We present NAO: a Nondeterministic tolerance Aware Optimistic verification protocol for floating-point neural networks that accepts outputs within principled operator-level acceptance regions rather than requiring bitwise equality. NAO combines two complementary error models: (i) sound per-operator IEEE-754 worst-case bounds and (ii) tight empirical percentile profiles calibrated across hardware. Discrepancies are resolved via a Merkle-anchored, threshold-guided interactive dispute game that recursively partitions the traced computation graph until one operator remains; at the leaf, adjudication reduces to either a lightweight theoretical-bound check or a small honest-majority vote against empirical thresholds. Unchallenged results finalize after a challenge window, without requiring trusted hardware or deterministic kernels.
We implement NAO as a PyTorch-compatible runtime and a contract layer currently deployed on Ethereum Holesky testnet. The runtime instruments graphs, computes per-operator bounds in real time, and runs unmodified vendor kernels in FP32 with negligible overhead (0.3% additional latency on Qwen3-8B). Our evaluation spans CNNs, Transformers, LLMs, and diffusion models across multiple GPUs (A100, H100, RTX6000, RTX4090). Results show that empirical thresholds are $10^{2}$$10^{3}\times$ tighter than theoretical bounds, and bound-aware adversarial attacks achieve 0% success under empirical checks. Together, NAO reconciles scalability with verifiability for real-world heterogeneous ML compute.
@@ 1. Introduction
Modern machine learning (ML) systems increasingly outsource training and inference to third-party infrastructure, including public clouds, inference marketplaces, edge clusters, and specialized accelerators (Zhang et al., 2019). This outsourcing is driven by the rapidly growing cost of ML computation, which has led to two notable trends. First, to conserve scarce compute budgets, practitioners often trade accuracy for cost at serving time, for example, by switching to smaller models, lowering precision, or coarsening quality knobs. Second, platform opacity makes these trade-offs invisible: in ML-as-a-service (MLaaS) platforms, whether closed-source proprietary models or open-model inference services, users cannot see which model was executed, what inputs were processed, or whether results faithfully reflect the intended computation. As a result, users accept outputs they cannot independently verify, leaving them exposed to silent downgrades such as model swaps, early exit (Lahiany and Aperstein, 2022), or quantization (Zhao et al., 2024), or service discrepancies like altered advertisement embeddings. When consumers of results do not control the underlying hardware, verifiable ML becomes a crucial system problem: the user should be provided with an efficiently verifiable proof certifying that the output was generated by executing an agreed model on agreed inputs (Jia et al., 2021; Choi et al., 2023).
@@
As demand for ML compute grows, platforms with decentralized inference markets and heterogeneous GPUs fleets promise greater capacity and price diversity. However, they also expose a fundamental technical obstacle for verifiable ML: modern accelerators do not guarantee deterministic outputs. Floating-point arithmetic (IEEE-754) (Kahan, 1996) is non-associative; vendor kernels legitimately reorder reductions and fuse operators; thread scheduling and atomics introduce nondeterministic execution traces (Chou et al., 2020; Shanmugavelu et al., 2024). Consequently, running the same neural network (NN) on different hardware, or even repeated executions on the same GPU can yield slightly different outputs. Here, we use operator to denote a primitive tensor-level function (a node in the NN graph), such as matrix multiplication, convolution, element-wise addition, or activation in PyTorch/ATen. In short, cross-platform nondeterminism is intrinsic to production GPU stacks, yet most existing verification methods assume exact reproducibility. This work addresses the gap.
Figure 1. Tolerance-aware optimistic verification on two example tasks. Left: Using LLM agents for automated transaction Honest proposer ($y_{0}$) returns the correct destination: 0xA173****d422, amount: 0.01. Malicious ($y_{1}$) applies unbounded node-wise perturbations and outputs an absurd amount to send (rejected). ($y_{2}$) stays within IEEE-754 worst-case bounds but violates empirical thresholds (slashed). ($y_{3}$) fits both bounds (accepted). Right: Image classification the honest proposer outputs “Cat”, while three similar types of malicious proposers output “Rock” (rejected), “Rabbit” (rejected), and “Cat” (accepted). Middle: tolerance-aware check. Malicious proposers inject perturbations after each operator. Results are rejected if they fall outside IEEE-754 theoretical bounds or empirical thresholds, while bounded, no explicit deviations are accepted. ($\oplus$: post-operator add)
@@
Why existing approaches fall short. (1) zkML: A natural path is to prove every arithmetic operation in zero knowledge. However, current end-to-end zkML pipelines remain orders of magnitude slower and memory-intensive for realistic floating point models (Wang and Hoang, 2023; Chen et al., 2024; Qu et al., 2025; Sun et al., 2024). Expressing large PyTorch graphs in field arithmetic or fixed-point circuits, and proving native FP32 GPU kernels at scale, is beyond today’s practicality. Moreover, real GPU stacks are inherently nondeterministic: one must either fix a specific execution schedule (sacrificing performance and hardware heterogeneity) or encode acceptance intervals, neither of which current zkML frameworks support. (2) Deterministic replay: Another approach is to enforce deterministic libraries, kernels, and reduction orders so that independent re-executions are bitwise reproducible (Zheng et al., 2021). This discards mature vendor libraries, requires extensive manual redesign of parallel reductions, and constrains accelerators (GPUs/TPUs) that are fundamentally nondeterministic by design. (3) Trusted execution environments (TEEs) attest that a computation ran inside an trusted enclave (Tramer and Boneh, [n. d.]; Moon et al., 2025; Bai et al., 2025). But this shifts trust to single-vendor microcode, introduces side-channel risks (Wei et al., 2018) and performance degradation (Niu et al., 2022) due to limited storage and compute capacity. (4) Replication and majority voting: Protocol-level schemes replicate the computation across multiple providers and accept the majority output. The high cost of replication for large models severely limits practicality and scalability (Zhao et al., 2005; Esmaeili and Mtibaa, 2024). In summary, none of these approaches simultaneously accommodate floating-point nondeterminism, preserve production-level performance, and enable portable, vendor-agnostic verification.
@@
This paper: NAO. We present NAO, an tolerance-aware optimistic execution protocol for verifiable neural network computation on heterogeneous accelerators. Instead of requiring bitwise equality, NAO verifies results up to model-specific, operator-level acceptance regions-sets of outputs guaranteed to remain within a prescribed error bound. Discrepancies are resolved through a Merkle-anchored interactive dispute game that localizes disagreements to a single operator (a leaf operator in the Merkle tree (Merkle, 1987)). At that leaf, lightweight adjudication suffices, either through a certified theoretical bound check or a small committee vote against pre-calibrated empirical thresholds. The key idea is to embrace, rather than eliminate, floating-point nondeterminism by verifying neural outputs up to principled per-operator error bounds, while preserving native GPU performance and hardware heterogeneity. By pushing verification down to the granularity where it is both meaningful and efficient: NAO reconciles scalability with accountability. Users can run the models they want, on any hardware the market provides, without compromising correctness, transparency, or recourse.
Two complementary error models. Even when software determinism is enforced (e.g., fixed kernel choices, disabled autotuning), parallel execution still reorders operations, introducing nondeterminism due to the non-associativity of IEEE-754 arithmetic. Rather than attempting to eliminate this through brittle execution constraints, we model and bound rounding error directly. NAO introduces two complementary error models that make tolerance-aware verification both principled and deployable:
@@
Our traced runtime instruments PyTorch graphs (Ansel et al., 2024), computes theoretical bounds on the fly, and records intermediate results to build robust cross-device empirical percentile profiles. Instead of propagating worst-case error across the entire graph, we turn composition into localization: disagreements are recursively narrowed until a single operator can be certified. By combining both error bounds, NAO flexibly balances universality, tightness, and scalability.
@@
From whole-model proofs to single-operator checks. Verifying a large model end-to-end against any bound remains expensive. NAO makes verification cheap by localizing disagreement and reducing the verification burden. Model owners provide cryptographic commitments to the weights and graph topology. In the common case, a proposer (compute provider) runs the model and posts a commitment; if unchallenged within a window, the commitment attains finality. If challenged, an interactive $N$-way partition protocol is invoked. At each round, empirical thresholds guide the split: the proposer partitions the traced graph into disjoint subgraphs with inclusion proofs, and the challenger identifies the first child whose outputs exceed its empirical thresholds. After $O(\log_{N}|V|)$ rounds ($V$ is the set of operators), the interactive dispute game narrows to a single operator with agreed inputs, where a theoretical-bound check or a committee vote evaluation resolves the dispute.
@@
Coordination layer. The protocol needs a coordination service that (a) stores tamper-evident commitments to model, graph, and outputs; (b) manages challenge windows and state transitions; and (c) escrows bonds and enforces rewards and punishments. One straightforward option is to deploy a trusted centralized coordinator. In our prototype, however, we instantiate this layer on a decentralized ledger (Ethereum), which provides authenticated logs, fair timeouts, and bond management out of the box, while also offering a transparent, widely understood cost model. Importantly, NAO itself does not rely on blockchain-specific assumptions.
@@
Open-model setting and service transparency. NAO targets the prevalent and rapidly growing open-model setting, where both weights and graph structure are public (e.g., community LLMs). This transparency enables subgraph extraction, per-operator thresholding, and provenance commitments. Closed-model APIs can also benefit, if providers expose such commitments to permissioned parties.
@@
Adversary and attack coverage. Tolerance-aware verification raises two concerns: trustworthiness and attack surface given the error bounds. We adopt a strong threat model in which the proposer may be malicious, injecting perturbations into intermediate tensors to evade empirical thresholds or theoretical error bound checks. To evaluate this risk, we design adaptive gradient-based attacks (Madry et al., 2018): distributional evasion against empirical envelopes and element-wise evasion against theoretical bounds. We further study whether many small, locally admissible deviations could accumulate into a task-level error without being detected. Our results show that the admissible per-operator perturbations are small and do not naturally compound, making end-to-end attacks under NAO practically difficult. Example is in Fig. 1.
@@
Implementation overview. We implement NAO as a PyTorch runtime and a set of Ethereum smart contracts. The PyTorch runtime instruments computation graphs, computes per-operator bounds on the fly, records intermediate traces, partitions and re-executes subgraphs, and emits or verifies Merkle commitments. The smart contracts manage commitments, dispute state transitions, per-round bonds, and slashing. The runtime leverages software-level determinism features, runs unmodified vendor kernels in FP32, and uses FP64 arithmetic for error-bound calculations. It introduces negligible overhead in the optimistic execution ($\approx$0.3% additional latency on Qwen3-8B) and requires no extra memory beyond native subgraph execution (Sec. 7).
@@
    Error analysis and adversarial attacks at operator granularity. We develop dual error models: portable theoretical IEEE-754 bounds and tight empirical percentile profiles, and conduct a comprehensive study of bound-aware adversarial attacks, quantifying their limited ability to subvert NAO across representative models.
@@
    Deployable PyTorch runtime and end-to-end system. We implement a PyTorch-compatible runtime that instruments models to compute and apply these error models in real time, integrate it into the full NAO pipeline, and demonstrate deployability in the open-model setting.
@@
Outline of the paper. We first present our system architecture and protocol overview in Sec. 3, followed by detailed methodology of two error models in Sec. 4, the adversary model and attack experiments in Sec. 5. We introduce the optimistic dispute game in Sec. 6 and evaluate our system’s performance, security, and cost in Sec. 7. Related work is in Sec. 8 with discussion and conclusion in Sec. 9.
2. Background
2.1. Nondeterminism is the default (replication is brittle)
In production, floating-point execution order shifts with GPU count/type and library stacks (cuBLAS/cuDNN/ATen), drivers, and runtimes. Bitwise replay occurs only when all knobs match: a condition that is impractical at scale and in decentralized markets. Forcing determinism often sacrifices performance and hardware heterogeneity. We empirically verified that endpoint settings meant to reduce randomness do not guarantee identical outputs across invocations. Concretely, on the OpenAI API we issued the same prompt to the same model with temperature T=0T{=}0, nucleus parameter top-p=1p{=}1, number of candidates n=1n{=}1, and a fixed seed $s$. Despite controlling both input randomness and output randomness (sampler), we observed different token sequences on repeated calls; when requesting token-level log probabilities, the reported logprobs also differed across runs. Similar caveats apply to other hosted services: Amazon Bedrock’s documentation states responses can be made more deterministic but not fully deterministic. [^1] https://repost.aws/knowledge-center/bedrock-prompts-deterministic-responses These observations highlight a core constraint for verifiers: exact replay at the service boundary is brittle and costly, even before accounting for heterogeneous hardware.
2.2. Benign numerical drift is ubiquitous and regular
NAO takes a different path: instead of suppressing nondeterminism, we model the numerical variance caused by benign reorderings. We measure post-operator runtime error, i.e., the element-wise deviation between two valid executions of the same operator given the same inputs but on different (hardware, library) configurations. Using traced models, we ran the identical operator sequences on two GPU types (H100 and RTX6000) and recorded the intermediate operator outputs along the graph’s canonical topological order. Fig. 2 shows that per-op deviations are small, largely flat across depth, and feature only localized spikes. Per-operator error histograms (not shown) are sharply peaked with light tails. This regularity suggests an opportunity: tolerance-aware verification that accepts outputs within principled, op-specific acceptance regions without fixing a global execution order.

### v3

# TAO: Tolerance-Aware Optimistic Verification for Floating-Point Neural Networks

@@ Abstract.
Neural networks increasingly run on hardware outside the user’s control (cloud GPUs, inference marketplaces, edge specialized accelerators) for both training and inference. Yet ML-as-a-Service reveals little about what actually ran or whether returned outputs faithfully reflect the intended inputs and models. Users lack recourse against service downgrades such as model swaps, quantization, graph rewrites, or discrepancies like altered advertisement embeddings. Verifying outputs is especially difficult because floating-point execution on heterogeneous accelerators is inherently nondeterministic. Existing approaches like zkML, deterministic replay, TEEs, and replication are either impractical for real floating-point neural networks or reintroduce vendor trust. We present TAO: a Tolerance Aware Optimistic verification protocol for floating-point neural networks that accepts outputs within principled operator-level acceptance regions rather than requiring bitwise equality. TAO combines two complementary error models: (i) sound per-operator IEEE-754 worst-case bounds and (ii) tight empirical percentile profiles calibrated across hardware. Discrepancies are resolved via a Merkle-anchored, threshold-guided interactive dispute game that recursively partitions the traced computation graph until one operator remains; at the leaf, adjudication reduces to either a lightweight theoretical-bound check or a small honest-majority vote against empirical thresholds. Unchallenged results finalize after a challenge window, without requiring trusted hardware or deterministic kernels.
We implement TAO as a PyTorch-compatible runtime and a coordinator contract layer; in our prototype, the coordinator is instantiated as an Ethereum smart-contract deployment (Holesky testnet) to quantify coordination cost. The runtime instruments graphs, computes per-operator bounds in real time, and runs unmodified vendor kernels in FP32 with negligible overhead (0.3% additional latency on Qwen3-8B). Our evaluation spans CNNs, Transformers, LLMs, and diffusion models across multiple GPUs (A100, H100, RTX6000, RTX4090). Results show that empirical thresholds are $10^{2}$$10^{3}\times$ tighter than theoretical bounds, and bound-aware adversarial attacks achieve 0% success under empirical checks. Together, TAO reconciles scalability with verifiability for real-world heterogeneous ML compute.
@@ 1. Introduction
Modern machine learning (ML) systems increasingly outsource training and inference to third-party infrastructure, including public clouds, inference marketplaces, edge clusters, and specialized accelerators (zhang2019empirical). This outsourcing is driven by the rapidly growing cost of ML computation, which has led to two notable trends. First, to conserve scarce compute budgets, practitioners often trade accuracy for cost at serving time, for example, by switching to smaller models, lowering precision, or coarsening quality knobs. Second, platform opacity makes these trade-offs invisible: in ML-as-a-service (MLaaS) platforms, whether closed-source proprietary models or open-model inference services, users cannot see which model was executed, what inputs were processed, or whether results faithfully reflect the intended computation. As a result, users accept outputs they cannot independently verify, leaving them exposed to silent downgrades such as model swaps, early exit (lahiany2022pteenet), or quantization (zhao2024atom), or service discrepancies like altered advertisement embeddings. When consumers of results do not control the underlying hardware, verifiable ML becomes a crucial system problem: the user should be provided with an efficiently verifiable proof certifying that the output was generated by executing an agreed model on agreed inputs (jia2021proof; choi2023tools).
Figure 1. Tolerance-aware optimistic verification for floating-point NN inference: the honest proposer outputs “Cat”; malicious proposers inject perturbations after each operator. Results are rejected if they fall outside IEEE-754 theoretical bounds or empirical thresholds, while bounded, no label-flip deviations are accepted.
@@
As demand for ML compute grows, platforms with decentralized inference markets and heterogeneous GPUs fleets promise greater capacity and price diversity. However, they also expose a fundamental technical obstacle for verifiable ML: modern accelerators do not guarantee deterministic outputs. Floating-point arithmetic (IEEE-754) (kahan1996ieee) is non-associative; vendor kernels legitimately reorder reductions and fuse operators; thread scheduling and atomics introduce nondeterministic execution traces (chou2020deterministic; shanmugavelu2024impacts). Consequently, running the same neural network (NN) on different hardware, or even repeated executions on the same GPU can yield slightly different outputs. Here, we use operator to denote a primitive tensor-level function (a node in the NN graph), such as matrix multiplication, convolution, element-wise addition, or activation in PyTorch/ATen. In short, cross-platform nondeterminism is intrinsic to production GPU stacks, yet most existing verification methods assume exact reproducibility. This work addresses the gap.
@@
Why existing approaches fall short. (1) zkML: A natural path is to prove every arithmetic operation in zero knowledge. However, current end-to-end zkML pipelines remain orders of magnitude slower and memory-intensive for realistic floating point models (wang2023ezdps; chen2024zkml; qu2025zkgpt; sun2024zkllm). Expressing large PyTorch graphs in field arithmetic or fixed-point circuits, and proving native FP32 GPU kernels at scale, is beyond today’s practicality. Moreover, real GPU stacks are inherently nondeterministic: one must either fix a specific execution schedule (sacrificing performance and hardware heterogeneity) or encode acceptance intervals, neither of which current zkML frameworks support. (2) Deterministic replay: Another approach is to enforce deterministic libraries, kernels, and reduction orders so that independent re-executions are bitwise reproducible (zheng2021agatha). This discards mature vendor libraries, requires extensive manual redesign of parallel reductions, and constrains accelerators (GPUs/TPUs) that are fundamentally nondeterministic by design. (3) Trusted execution environments (TEEs) attest that a computation ran inside an trusted enclave (tramerslalom; moon2025asgard; bai2025phantom). But this shifts trust to single-vendor microcode, introduces side-channel risks (wei2018know) and performance degradation (niu20223legrace) due to limited storage and compute capacity. (4) Replication and majority voting: Protocol-level schemes replicate the computation across multiple providers and accept the majority output. The high cost of replication for large models severely limits practicality and scalability (zhao2005result; esmaeili2024serene). In summary, none of these approaches simultaneously accommodate floating-point nondeterminism, preserve production-level performance, and enable portable, vendor-agnostic verification.
@@
This paper: TAO. We present TAO, an tolerance-aware optimistic execution protocol for verifiable neural network computation on heterogeneous accelerators. Instead of requiring bitwise equality, TAO verifies results up to model-specific, operator-level acceptance regions-sets of outputs guaranteed to remain within a prescribed error bound. Discrepancies are resolved through a Merkle-anchored interactive dispute game that localizes disagreements to a single operator (a leaf operator in the Merkle tree (merkle1987digital)). At that leaf, lightweight adjudication suffices, either through a certified theoretical bound check or a small committee vote against pre-calibrated empirical thresholds. The key idea is to embrace, rather than eliminate, floating-point nondeterminism by verifying neural outputs up to principled per-operator error bounds, while preserving native GPU performance and hardware heterogeneity. By pushing verification down to the granularity where it is both meaningful and efficient: TAO reconciles scalability with accountability. Users can run the models they want, on any hardware the market provides, without compromising correctness, transparency, or recourse.
Two complementary error models. Even when software determinism is enforced (e.g., fixed kernel choices, disabled autotuning), parallel execution still reorders operations, introducing nondeterminism due to the non-associativity of IEEE-754 arithmetic. Rather than attempting to eliminate this through brittle execution constraints, we model and bound rounding error directly. TAO introduces two complementary error models that make tolerance-aware verification both principled and deployable:
@@
Our traced runtime instruments PyTorch graphs (ansel2024pytorch), computes theoretical bounds on the fly, and records intermediate results to build robust cross-device empirical percentile profiles. Instead of propagating worst-case error across the entire graph, we turn composition into localization: disagreements are recursively narrowed until a single operator can be certified. By combining both error bounds, TAO flexibly balances universality, tightness, and scalability.
@@
From whole-model proofs to single-operator checks. Verifying a large model end-to-end against any bound remains expensive. TAO makes verification cheap by localizing disagreement and reducing the verification burden. Model owners provide cryptographic commitments to the weights and graph topology. In the common case, a proposer (compute provider) runs the model and posts a commitment; if unchallenged within a window, the commitment attains finality. If challenged, an interactive $N$-way partition protocol is invoked. At each round, empirical thresholds guide the split: the proposer partitions the traced graph into disjoint subgraphs with inclusion proofs, and the challenger identifies the first child whose outputs exceed its empirical thresholds. After $O(\log_{N}|V|)$ rounds ($V$ is the set of operators), the interactive dispute game narrows to a single operator with agreed inputs, where a theoretical-bound check or a committee vote evaluation resolves the dispute.
@@
Coordination layer. The protocol needs a coordination service that (a) stores tamper-evident commitments to model, graph, and outputs; (b) manages challenge windows and state transitions; and (c) escrows bonds and enforces rewards and punishments. One straightforward option is to deploy a trusted centralized coordinator. In our prototype, however, we instantiate this layer on a decentralized ledger (Ethereum), which provides authenticated logs, fair timeouts, and bond management out of the box, while also offering a transparent, widely understood cost model. Importantly, TAO itself does not rely on blockchain-specific assumptions.
@@
Open-model setting and service transparency. TAO targets the prevalent and rapidly growing open-model setting, where both weights and graph structure are public (e.g., community LLMs). This transparency enables subgraph extraction, per-operator thresholding, and provenance commitments. Closed-model APIs can also benefit, if providers expose such commitments to permissioned parties.
@@
Adversary and attack coverage. Tolerance-aware verification raises two concerns: trustworthiness and attack surface given the error bounds. We adopt a strong threat model in which the proposer may be malicious, injecting perturbations into intermediate tensors to evade empirical thresholds or theoretical error bound checks. To evaluate this risk, we design adaptive gradient-based attacks (madry2018towards): distributional evasion against empirical envelopes and element-wise evasion against theoretical bounds. We further study whether many small, locally admissible deviations could accumulate into a task-level error without being detected. Our results show that the admissible per-operator perturbations are small and do not naturally compound, making end-to-end attacks under TAO practically difficult. Example is in Fig. 1.
@@
Implementation overview. We implement TAO as a PyTorch runtime and a set of Ethereum smart contracts as the coordinator. The PyTorch runtime instruments computation graphs, computes per-operator bounds on the fly, records intermediate traces, partitions and re-executes subgraphs, and emits or verifies Merkle commitments. The smart contracts manage commitments, dispute state transitions, per-round bonds, and misbehavior penalties. The runtime leverages software-level determinism features, runs unmodified vendor kernels in FP32, and uses FP64 arithmetic for error-bound calculations. It introduces negligible overhead in the optimistic execution ($\approx$0.3% additional latency on Qwen3-8B) and requires no extra memory beyond native subgraph execution (Sec. 6).
@@
    Error analysis and adversarial attacks at operator granularity. We develop dual error models: portable theoretical IEEE-754 bounds and tight empirical percentile profiles, and conduct a comprehensive study of bound-aware adversarial attacks, quantifying their limited ability to subvert TAO across representative models.
@@
    Deployable PyTorch runtime and end-to-end system. We implement a PyTorch-compatible runtime that instruments models to compute and apply these error models in real time, integrate it into the full TAO pipeline, and demonstrate deployability in the open-model setting.
@@
Outline of the paper. We first present our system architecture and protocol overview in Sec. 2, followed by detailed methodology of two error models in Sec. 3, the adversary model and attack experiments in Sec. 4. We introduce the optimistic dispute game in Sec. 5 and evaluate our system’s performance, security, and cost in Sec. 6. Related work is in Sec. 8 with discussion  Sec.7 and conclusion in Sec. 9.
2. System architecture and protocol overview